OpenStudy (anonymous):

plot the graph of the following equation for -3 *less than equal to sign here* x *less than equal to sign here* 3 Also a table needs to be filled out for this y=2^x+2

OpenStudy (ikram002p):

\(-3\le x \le 3\) also sketch this ? \(y=2^x+2 \) ??

OpenStudy (ikram002p):

or you wanna sketch \(y=2^x+2\) with range of \(-3\le x\le 3\)

OpenStudy (anonymous):

yes sketch y=2^x+2 with those ranges

OpenStudy (ikram002p):

well i guess is the second one simply fill this table :- 

OpenStudy (ikram002p):

ill solve first one , so lets see if you could continue with others ok ?

OpenStudy (anonymous):

ok thank you

OpenStudy (ikram002p):